Two lines through the point (1,3) are tangent to the curve y=x^2. find the equations of these two lines and make a sketch to verify your result
Answer by josgarithmetic(39617)

Where is point (1,3) in relation to the curve y=x^2?
y=1*1=2, when x=1 for the curve.

The point (1,3) is ABOVE the curve, a parabola opening upward. The given point is NOT on the part of the plane necessary for being part of any line tangent to the parabola.

Your question has no solution.